A bank pays 6% interest per year, compounded quarterly. To what amount will a $5000 deposit grow if left in that bank for 10 years? Ans. $9070.09
Answer #1

Future value = Present value (1+R)^N

= 5000 (1+0.06/4)^10*4

=5000 (1+0.015)*40

=5000* 1.8140 i.e 9070.09